# Gauge-Invariant Longitudinal Modes in the Herwig 7 Electroweak Parton Shower

### v1

@@ Abstract
Longitudinal electroweak gauge bosons are the most technically delicate ingredient of electroweak parton showers: in the broken Standard Model, the gauge component of a longitudinal polarisation does not cancel diagram by diagram, but is related by Ward identities to amplitudes with an insertion of the associated would-be Goldstone field. Building on the default Herwig 7 treatment based on the Dawson-subtracted longitudinal current, we construct a gauge-invariant longitudinal scheme in which the Dawson remainder is retained as a well-defined contribution and completed by a Ward-identity-fixed Goldstone-matching term. We derive helicity-resolved building blocks and compact quasi-collinear splitting kernels for $q\to q^{\prime}V$ and $V\to V^{\prime}V^{\prime\prime}$ branchings, and implement the scheme in Herwig 7 as a switchable alternative to the default longitudinal basis. The completion leaves the transverse sector unchanged and modifies only longitudinal entries through controlled symmetry-breaking terms, including Yukawa-sensitive contributions in massive-fermion channels. In shower-level studies, we find that the Dawson and gauge-invariant prescriptions coincide at high evolution scales, while differing at lower scales precisely in channels where symmetry breaking is active. Exclusive single-emission observables can moreover display non-monotonic scheme dependence once Sudakov suppression and kinematic constraints are accounted for. A first multi-emission LHC-like study confirms numerical stability and yields controlled, interpretable shifts in observables that probe propagating electroweak shower currents, whereas quantities dominated by promptly contracted, near on-shell vector-boson matrix elements remain largely unchanged.
@@ 1 Introduction
Electroweak (EW) parton showers have become an essential component of precision and discovery phenomenology at present and future hadron colliders, because in the multi-TeV regime the exclusive structure of events is increasingly shaped by real and virtual EW radiation and by the associated Sudakov-enhanced logarithms, schematically $\alpha_{\mathrm{w}}\ln^{2}(Q^{2}/M_{W}^{2})$ and $\alpha_{\mathrm{w}}\ln(Q^{2}/M_{W}^{2})$, which can compete numerically with QCD effects in energetic tails and are not reliably captured by fixed-order calculations alone once widely separated scales and multiple emissions are relevant Ciafaloni:2000df ; Denner:2000jv ; Kuhn:1999nn . Implementing this resummation in a fully differential Monte Carlo framework, however, is conceptually more subtle than in QCD, because the broken gauge theory introduces physical masses, chiral couplings, and mixing, and in particular renders the treatment of longitudinally polarised massive vector bosons delicate: the conventional polarisation vector $\varepsilon_{L}^{\mu}(p)$ contains a formally large $p^{\mu}/M_{V}$ component whose contributions are gauge artefacts that cancel only after enforcing the appropriate Ward identities, with the cancellation controlled, at leading power in $M_{V}^{2}/Q^{2}$, by the Goldstone-boson equivalence theorem and by Yukawa interactions for massive fermions Cornwall:1974km ; Chanowitz:1985hj ; Denner:2001gw . In a shower context, where one factorises a hard process from quasi-collinear branchings and interprets the latter probabilistically, any ad hoc subtraction or redefinition of $\varepsilon_{L}^{\mu}$ that removes the $p^{\mu}/M_{V}$ term must therefore be accompanied by a correspondingly consistent inclusion of the would-be Goldstone contribution; otherwise one risks violating gauge invariance at the level of the splitting operator and, practically, suppressing or mis-modelling longitudinal configurations precisely in the ultracollinear region (with transverse momenta of order $M_{V}$, where mass effects are parametrically important) Masouminia:2021kne ; Dittmaier:2025htf . The state-of-the-art consequently combines helicity-resolved (or spin-density-matrix) evolution with quasi-collinear EW splitting kernels in the broken theory, together with prescriptions, either in physical gauges or via explicit Goldstone-matching, that maintain gauge-invariant (GI) longitudinal currents at the level of the splitting operator while retaining the correct on-shell/virtual interpolation for $W/Z/H$ emissions and preserving spin correlations to the accuracy claimed by the shower Masouminia:2021kne ; Christiansen:2014kba ; Krauss:2014yaa ; Bauer:2017isx .
@@
In Herwig 7 Bahr:2008pv ; Bellm:2015jjp ; Bellm:2017bvx ; Bellm:2019zci ; Bewick:2023tfi ; Bellm:2025pcw , the EW parton shower was formulated as an extension of the angular-ordered, spin-correlated shower to the spontaneously broken $\mathrm{SU}(2)_{L}\times\mathrm{U}(1)_{Y}$ theory, with quasi-collinear splitting operators for fermion, vector-boson and scalar lines embedded into the evolution via spin-density matrices in order to retain leading spin correlations within a Markovian branching framework Masouminia:2021kne ; Darvishi:2021het ; Bellm:2025pcw . The longitudinal sector was handled by adopting the “Dawson picture”: longitudinal kernels were constructed using a manifestly finite definition of the longitudinal state which avoids the explicit appearance of the $p^{\mu}/M_{V}$ component of $\varepsilon_{L}^{\mu}(p)$ at the level of the branching operator, thereby providing numerically stable helicity-dependent splitting functions suitable for shower evolution Masouminia:2021kne . This prescription is well-motivated in Equivalent Vector Boson Approximation (EWA)-like kinematics, and fixed-order comparisons in that regime provided non-trivial validation of the resulting resummation pattern Masouminia:2021kne ; nevertheless, recent work has emphasised that, outside this restricted limit, the structure of longitudinal contributions in the broken theory is constrained by Ward identities which relate the would-be $p^{\mu}$ contraction to amplitudes involving the corresponding Goldstone mode (and, for massive fermions, Yukawa interactions), raising the question of whether a universal application of the Dawson construction across all EW branchings can miss parts of the ultracollinear longitudinal dynamics Dittmaier:2025htf . The resulting discussion is therefore not about the practicality of the Dawson implementation, but about its domain of formal validity and the extent to which a GI completion is required to control longitudinal rates and polarisation systematics in general shower applications.
@@
Conceptually, it is useful to distinguish between the Dawson picture and a fully GI treatment of longitudinal modes. In the Dawson picture, one replaces the conventional longitudinal polarisation vector by a subtracted object, $\varepsilon_{L}^{\mu}(p)\to\varepsilon_{L*}^{\mu}(p)$, obtained by removing the component proportional to $p^{\mu}/M_{V}$, so that the splitting currents entering the quasi-collinear kernels are manifestly finite and directly amenable to a probabilistic shower interpretation Masouminia:2021kne ; this construction is particularly natural in EWA-like kinematics, where the subtracted $p^{\mu}/M_{V}$ contribution is power-suppressed or eliminated once the appropriate Ward identities are enforced on the hard amplitude Masouminia:2021kne . In the spontaneously broken theory, however, the relevant Ward identities do not, in general, imply that contractions with $p^{\mu}$ vanish: instead, they relate them to amplitudes with an insertion of the associated would-be Goldstone field (with Yukawa-suppressed contributions for massive fermions), so that the physically meaningful longitudinal limit is governed by a specific vector-Goldstone combination Cornwall:1974km ; Chanowitz:1985hj ; Denner:2001gw . A GI longitudinal scheme therefore cannot be realised by subtraction alone; rather, it promotes the Dawson remainder to one component of the physical longitudinal current and supplements it by the Ward-identity-fixed Goldstone-matching term (for our conventions), yielding a longitudinal splitting operator which is gauge invariant by construction and applicable to both on-shell and off-shell shower kinematics across the full set of EW branching channels Dittmaier:2025htf .
@@
In this paper, we formulate and implement such a GI completion of the longitudinal sector of the Herwig 7 electroweak parton shower. We derive the corresponding helicity-resolved building blocks and compact quasi-collinear splitting kernels for both fermion and vector-boson branchings, and implement the scheme in Herwig 7 in a manner which preserves backwards compatibility and allows a transparent comparison between the Dawson and GI pictures within an otherwise identical shower setup. We validate the construction by analytic checks of the relevant Ward identities and by numerical comparisons at the level of splitting functions and controlled shower evolutions, isolating the regions of phase space in which the GI completion affects longitudinal rates and kinematics; the formal construction is presented in Sec. 3.
@@
The remainder of this paper is organised as follows. In Sec. 2 we summarise the Herwig 7 conventions for shower kinematics and evolution variables relevant for the EW shower, including the dot-product preserving reconstruction used throughout our validation studies. In Sec. 3 we develop the GI treatment of longitudinal polarisations in the broken theory, formulating a longitudinal splitting operator consistent with the relevant Ward identities and the associated Goldstone-matching, and specifying its relation to the Dawson picture used previously. Our results are presented in Sec. 4: we first compare the Dawson and GI schemes at the level of helicity building blocks and analytic quasi-collinear splitting functions, and then study their impact within controlled shower evolutions, ranging from single-resummed tests to a first multi-emission LHC-like check, with an emphasis on longitudinally sensitive observables and polarisation systematics. We conclude in Sec. 5. Details of the Herwig user interface and steering of the longitudinal scheme are collected in App. A.
@@ 3 Quasi-Collinear Helicity-Dependent Splitting Functions
with $m_{0}$ the parent pole mass. The longitudinal vector $\epsilon^{\mu}_{0}(p)$ contains the familiar $p^{\mu}/m_{0}$ behaviour and is therefore the locus of the longitudinal-mode subtleties in the broken theory. The Herwig 7 EW shower, as originally formulated, instead employs a manifestly finite “Dawson” longitudinal remainder, obtained by subtracting the would-be gauge piece and which in this frame can be written as Masouminia:2021kne
@@
which reduces to the familiar unpolarised splitting function upon averaging $\rho_{0}$ over the parent helicities. The decomposition of $F_{0\to 1\,2}$ into transverse-transverse, longitudinal-longitudinal, and mixed transverse-longitudinal components is thus entirely controlled by the choice of longitudinal basis used to define $H$, and this is precisely where the distinction between the Dawson prescription and a GI longitudinal construction enters. In the Dawson picture, one effectively defines the longitudinal contribution by the replacement $\epsilon_{0}^{\mu}\to\epsilon_{0_{*}}^{\mu}$ at the level of the splitting current, thereby yielding a finite longitudinal kernel that is straightforwardly embedded into the shower Masouminia:2021kne . In a fully GI treatment, by contrast, the longitudinal splitting current is defined such that it obeys the broken-theory Ward identities, which relate contractions of the vector current with the boson momentum to amplitudes with insertions of the corresponding would-be Goldstone field (and Yukawa terms for massive fermions), implying that the physically meaningful longitudinal object is a specific vector-Goldstone combination rather than the Dawson remainder alone Borel:2012by ; Cuomo:2019siu ; Dittmaier:2025htf . The following two subsections make this statement concrete for $q\to q^{\prime}V$ and $V\to V^{\prime}V^{\prime\prime}$ respectively, and provide the explicit Dawson and GI definitions of the reduced helicity amplitudes $H$ from which the shower kernels are constructed. Splittings involving a physical Higgs, $q\to q^{\prime}S$ and $V\to V^{\prime}S$ ($S$ being any possible scalar boson), do not require a Dawson/Goldstone completion of a longitudinal current in the sense discussed above, and their phenomenology is controlled by the corresponding Yukawa- and mass-suppressed structures.
@@ 3.1 $q\to q^{\prime}V$ splittings
with $\mathrm{X}=\mathrm{Dawson},\mathrm{GI}$ denoting the longitudinal prescription and with $\varepsilon_{\mu,\lambda_{2}}$ evaluated using the kinematics and phase conventions of Sec. 2. In Herwig 7, the corresponding splitting kernel is obtained by contracting $\widehat{H}$ with the fermion spin-density matrix $\rho$ of the emitting line and (optionally) retaining a spin-density matrix for the emitted vector for subsequent evolution/decay,
@@
where $0_{*}$ denotes the longitudinal state used in the shower (i.e. the Dawson-subtracted longitudinal current, or its GI completion), and where $F_{\lambda_{2}\lambda_{2}}$ stands for the corresponding helicity-diagonal contribution after summing $\lambda_{0},\lambda_{0}^{\prime},\lambda_{1}$ with $\rho$. The relevant Ward identity implies, schematically,
@@
up to our phase and coupling conventions. Accordingly, the GI longitudinal splitting current is obtained by supplementing the Dawson remainder by the corresponding Goldstone-matching term implied by (28).
Table 1 summarises the non-vanishing reduced amplitudes in the notation of Ref. Masouminia:2021kne , with the Dawson prescription corresponding to the pure subtracted longitudinal current and the GI prescription obtained by adding the uniquely fixed Goldstone-matching contribution implied by the broken-theory Ward identity (cf. Sec. 3). Concretely, the GI completion modifies only the longitudinal entries ($\lambda_{2}=0_{*}$), while leaving the transverse helicity amplitudes unchanged; the additional terms are proportional to the Goldstone-fermion couplings
@@
and a channel-dependent coefficient $c_{G}$ (encoding the normalisation/convention for the Goldstone insertion in the matching). At the amplitude-level, the GI completion can be formulated as:
$\widehat{H}^{(\mathrm{GI})}_{\lambda_{0}\lambda_{1}0_{*}}=\widehat{H}^{(\mathrm{Dawson})}_{\lambda_{0}\lambda_{1}0_{*}}+c_{G}\widehat{H}^{(\mathrm{G_{V}})}_{\lambda_{0}\lambda_{1}0_{*}}$, (30)
with $\widehat{H}^{(\mathrm{G_{V}})}_{\lambda_{0}\lambda_{1}}$ built from the scalar (Goldstone) Yukawa current, i.e. an Eq. (29)-type operator.
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{Dawson}}_{\lambda_{0}\lambda_{1}\lambda_{2}}=\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{Dawson}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$ $\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
Table 1: Non-vanishing reduced helicity amplitudes for $q\to q^{\prime}V$ in the Dawson and GI prescriptions. Entries common to both prescriptions are grouped in the upper block, while the lower block collects the longitudinal ($0_{*}$) components, for which the GI completion adds Goldstone-matching terms proportional to $c_{G}$ and $(y_{L},y_{R})$.
Using the decomposition $F^{(\mathrm{X})}=F_{T}+F_{L}^{(\mathrm{X})}$ and performing the helicity sums implied by the shower construction, weighted by the diagonal entries of the spin-density matrix $\rho$, one obtains the following closed forms for the transverse contribution and for the longitudinal contribution in the Dawson and GI prescriptions. Writing the kernels as functions of $(z,\tilde{q}^{2})$ (with $\tilde{q}$ the evolution variable of Sec. 2 and with $p_{\perp}^{2}$ eliminated using Eq. (7)), the transverse part is common to both schemes,
@@
whereas the Dawson longitudinal contribution reads
$F_{L;\rm{Dawson}}^{q\to q^{\prime}V}={m_{2}^{2}\over\tilde{q}^{2}(1-z)^{3}}\,\big(g_{R}^{2}\;\rho_{++}+g_{L}^{2}\;\rho_{--}\big),$ (32)
@@
Equations (31)-(33) make explicit the structural point emphasised in Sec. 3: the Dawson and GI prescriptions share an identical transverse sector, while their longitudinal sectors differ by terms controlled by the Goldstone-matching, including both interference ($\propto c_{G}$) and pure Goldstone ($\propto c_{G}^{2}$) contributions, which are parametrically suppressed for light fermions in Standard Model applications but become essential whenever symmetry-breaking effects and Yukawa insertions compete with the naive $p^{\mu}/M_{V}$ subtraction.

### v2

@@ Abstract
Longitudinal electroweak gauge bosons are the most technically delicate ingredient of electroweak parton showers: in the broken Standard Model, the gauge component of a longitudinal polarisation does not cancel diagram by diagram, but is related by Ward identities to amplitudes with an insertion of the associated would-be Goldstone field. Building on the default Herwig 7 treatment based on a subtraction-defined longitudinal current, we construct a gauge-invariant scheme in which the subtraction remainder is retained as a well-defined contribution and completed by a Ward-identity-fixed Goldstone-matching term. We derive helicity-resolved building blocks and compact quasi-collinear splitting kernels for $q\to q^{\prime}V$ and $V\to V^{\prime}V^{\prime\prime}$ branchings, and implement the scheme in Herwig 7 as a switchable alternative to the default basis. The completion leaves the transverse sector unchanged and modifies only longitudinal entries through controlled symmetry-breaking terms, including Yukawa-sensitive contributions in massive-fermion channels. In shower-level studies, we find that the default and gauge-invariant prescriptions coincide at high evolution scales, while differing at lower scales precisely in channels where symmetry breaking is active. Exclusive single-emission observables can moreover display non-monotonic scheme dependence once Sudakov suppression and kinematic constraints are accounted for. A first multi-emission LHC-like study confirms numerical stability and yields controlled, interpretable shifts in observables that probe propagating electroweak shower currents, whereas quantities dominated by promptly contracted, near on-shell vector-boson matrix elements remain largely unchanged.
@@ 1 Introduction
Electroweak (EW) parton showers have become an essential component of precision and discovery phenomenology at present and future hadron colliders, because in the multi-TeV regime the exclusive structure of events is increasingly shaped by real and virtual EW radiation and by the associated Sudakov-enhanced logarithms, schematically $\alpha_{\mathrm{w}}\ln^{2}(Q^{2}/M_{W}^{2})$ and $\alpha_{\mathrm{w}}\ln(Q^{2}/M_{W}^{2})$, which can compete numerically with QCD effects in energetic tails and are not reliably captured by fixed-order calculations alone once widely separated scales and multiple emissions are relevant Ciafaloni:2000df ; Denner:2000jv ; Kuhn:1999nn . Implementing this resummation in a fully differential Monte Carlo framework, however, is more subtle than in QCD: the broken gauge theory introduces physical masses, chiral couplings, and mixing, and in particular renders the treatment of longitudinal massive vector bosons delicate. The conventional polarisation vector $\varepsilon_{L}^{\mu}(p)$ contains a formally large $p^{\mu}/M_{V}$ component whose contributions are gauge artefacts that cancel only after enforcing the appropriate Ward identities, with the cancellation controlled, at leading power in $M_{V}^{2}/Q^{2}$, by the Goldstone-boson equivalence theorem and by Yukawa interactions for massive fermions Cornwall:1974km ; Chanowitz:1985hj ; Denner:2001gw . In a shower context, where one factorises a hard process from quasi-collinear branchings and interprets the latter probabilistically, any ad hoc subtraction or redefinition of $\varepsilon_{L}^{\mu}$ that removes the $p^{\mu}/M_{V}$ term must therefore be accompanied by a correspondingly consistent inclusion of the would-be Goldstone contribution; otherwise one risks violating gauge invariance at the level of the splitting operator and, practically, suppressing or mis-modelling such configurations precisely in the ultracollinear region (with transverse momenta of order $M_{V}$, where mass effects are parametrically important) Masouminia:2021kne ; Dittmaier:2025htf . The state-of-the-art consequently combines helicity-resolved (or spin-density-matrix) evolution with quasi-collinear EW splitting kernels in the broken theory, together with prescriptions, either in physical gauges or via explicit Goldstone-matching, that maintain gauge-invariant (GI) longitudinal currents at the level of the splitting operator while retaining the correct on-shell/virtual interpolation for $W/Z/H$ emissions and preserving spin correlations to the accuracy claimed by the shower Masouminia:2021kne ; Christiansen:2014kba ; Krauss:2014yaa ; Bauer:2017isx .
@@
In Herwig 7 Bahr:2008pv ; Bellm:2015jjp ; Bellm:2017bvx ; Bellm:2019zci ; Bewick:2023tfi ; Bellm:2025pcw , the EW parton shower was formulated as an extension of the angular-ordered, spin-correlated shower to the spontaneously broken $\mathrm{SU}(2)_{L}\times\mathrm{U}(1)_{Y}$ theory, with quasi-collinear splitting operators for fermion, vector-boson and scalar lines embedded into the evolution via spin-density matrices in order to retain leading spin correlations within a Markovian branching framework Masouminia:2021kne ; Darvishi:2021het ; Bellm:2025pcw . The longitudinal sector was handled by adopting a subtraction-longitudinal (SL) picture (also known as Dawson’s approach Dawson:1984gx ): longitudinal kernels were constructed using a manifestly finite definition of the longitudinal state which avoids the explicit appearance of the $p^{\mu}/M_{V}$ component of $\varepsilon_{L}^{\mu}(p)$ at the level of the branching operator, thereby providing numerically stable helicity-dependent splitting functions suitable for shower evolution Masouminia:2021kne . This prescription is well-motivated in Equivalent Vector Boson Approximation (EWA)-like kinematics, and fixed-order comparisons in that regime provided non-trivial validation of the resulting resummation pattern Masouminia:2021kne ; nevertheless, recent work has emphasised that, outside this restricted limit, the structure of longitudinal contributions in the broken theory is constrained by Ward identities which relate the would-be $p^{\mu}$ contraction to amplitudes involving the corresponding Goldstone mode (and, for massive fermions, Yukawa interactions), raising the question of whether a universal application of the SL construction across all EW branchings can miss parts of the ultracollinear longitudinal dynamics Dittmaier:2025htf . The resulting discussion is therefore not about the practicality of the SL implementation, but about its domain of formal validity and the extent to which a GI completion is required to control longitudinal rates and polarisation systematics in general shower applications.
@@
Conceptually, it is useful to distinguish between the SL picture and a fully GI treatment of longitudinal modes. In the SL picture, one replaces the conventional polarisation vector by a subtracted object, $\varepsilon_{L}^{\mu}(p)\to\varepsilon_{L*}^{\mu}(p)$, obtained by removing the component proportional to $p^{\mu}/M_{V}$, so that the splitting currents entering the quasi-collinear kernels are manifestly finite and directly amenable to a probabilistic shower interpretation Masouminia:2021kne ; this construction is particularly natural in EWA-like kinematics, where the subtracted $p^{\mu}/M_{V}$ contribution is power-suppressed or eliminated once the appropriate Ward identities are enforced on the hard amplitude Masouminia:2021kne . In the spontaneously broken theory, however, the relevant Ward identities do not, in general, imply that contractions with $p^{\mu}$ vanish: instead, they relate them to amplitudes with an insertion of the associated would-be Goldstone field (with Yukawa-suppressed contributions for massive fermions), so that the physical limit is governed by a specific vector-Goldstone combination Cornwall:1974km ; Chanowitz:1985hj ; Denner:2001gw . A GI longitudinal scheme therefore cannot be realised by subtraction alone; rather, it promotes the subtraction remainder to one component of the physical longitudinal current and supplements it by the Ward-identity-fixed Goldstone-matching term (for our conventions), yielding a longitudinal splitting operator which is gauge invariant by construction and applicable to both on-shell and off-shell shower kinematics across the full set of EW branching channels Dittmaier:2025htf .
@@
In this paper, we formulate and implement such a GI completion of the longitudinal sector of the Herwig 7 electroweak parton shower. We derive the corresponding helicity-resolved building blocks and compact quasi-collinear splitting kernels for both fermion and vector-boson branchings, and implement the scheme in Herwig 7 in a manner which preserves backwards compatibility and allows a transparent comparison between the SL and GI pictures within an otherwise identical shower setup. We validate the construction by analytic checks of the relevant Ward identities and by numerical comparisons at the level of splitting functions and controlled shower evolutions, isolating the regions of phase space in which the GI completion affects longitudinal rates and kinematics; the formal construction is presented in Sec. 3.
@@
The remainder of this paper is organised as follows. In Sec. 2 we summarise the Herwig 7 conventions for shower kinematics and evolution variables relevant for the EW shower, including the dot-product preserving reconstruction used throughout our validation studies. In Sec. 3 we develop the GI treatment of longitudinal polarisations in the broken theory, formulating a longitudinal splitting operator consistent with the relevant Ward identities and the associated Goldstone-matching, and specifying its relation to the SL picture used previously. Our results are presented in Sec. 4: we first compare the SL and GI schemes at the level of helicity building blocks and analytic quasi-collinear splitting functions, and then study their impact within controlled shower evolutions, ranging from single-resummed tests to a first multi-emission LHC-like check, with an emphasis on longitudinally sensitive observables and polarisation systematics. We conclude in Sec. 5. Details of the Herwig user interface and steering of the longitudinal scheme are collected in App. A.
@@ 3 Quasi-Collinear Helicity-Dependent Splitting Functions
with $m_{0}$ the parent pole mass. The longitudinal vector $\epsilon^{\mu}_{0}(p)$ contains the familiar $p^{\mu}/m_{0}$ behaviour and is therefore the locus of the longitudinal-mode subtleties in the broken theory. The Herwig 7 EW shower, as originally formulated, instead employs a manifestly finite SL remainder, obtained by subtracting the would-be gauge piece and which in this frame can be written as Masouminia:2021kne
@@
which reduces to the familiar unpolarised splitting function upon averaging $\rho_{0}$ over the parent helicities. The decomposition of $F_{0\to 1\,2}$ into transverse-transverse, longitudinal-longitudinal, and mixed transverse-longitudinal components is thus entirely controlled by the choice of longitudinal basis used to define $H$, and this is precisely where the distinction between the SL prescription and a GI longitudinal construction enters. In the SL picture, one effectively defines the longitudinal contribution by the replacement $\epsilon_{0}^{\mu}\to\epsilon_{0_{*}}^{\mu}$ at the level of the splitting current, thereby yielding a finite longitudinal kernel that is straightforwardly embedded into the shower Masouminia:2021kne . In a fully GI treatment, by contrast, the longitudinal splitting current is defined such that it obeys the broken-theory Ward identities, which relate contractions of the vector current with the boson momentum to amplitudes with insertions of the corresponding would-be Goldstone field (and Yukawa terms for massive fermions), implying that the physically meaningful longitudinal object is a specific vector-Goldstone combination rather than the subtraction remainder alone Borel:2012by ; Cuomo:2019siu ; Dittmaier:2025htf . The following two subsections make this statement concrete for $q\to q^{\prime}V$ and $V\to V^{\prime}V^{\prime\prime}$ respectively, and provide the explicit SL and GI definitions of the reduced helicity amplitudes $H$ from which the shower kernels are constructed. Splittings involving a physical Higgs, $q\to q^{\prime}S$ and $V\to V^{\prime}S$ ($S$ being any possible scalar boson), do not require a SL/Goldstone completion of a longitudinal current in the sense discussed above, and their phenomenology is controlled by the corresponding Yukawa- and mass-suppressed structures.
@@ 3.1 $q\to q^{\prime}V$ splittings
with $\mathrm{X}=\mathrm{SL},\mathrm{GI}$ denoting the longitudinal prescription and with $\varepsilon_{\mu,\lambda_{2}}$ evaluated using the kinematics and phase conventions of Sec. 2. In Herwig 7, the corresponding splitting kernel is obtained by contracting $\widehat{H}$ with the fermion spin-density matrix $\rho$ of the emitting line and (optionally) retaining a spin-density matrix for the emitted vector for subsequent evolution/decay,
@@
where $0_{*}$ denotes the longitudinal state used in the shower (i.e. the subtracted longitudinal current, or its GI completion), and where $F_{\lambda_{2}\lambda_{2}}$ stands for the corresponding helicity-diagonal contribution after summing $\lambda_{0},\lambda_{0}^{\prime},\lambda_{1}$ with $\rho$. The relevant Ward identity implies, schematically,
@@
up to our phase and coupling conventions. Accordingly, the GI longitudinal splitting current is obtained by supplementing the subtraction remainder by the corresponding Goldstone-matching term implied by (28).
Table 1 summarises the non-vanishing reduced amplitudes in the notation of Ref. Masouminia:2021kne , with the SL prescription corresponding to the pure subtracted longitudinal current and the GI prescription obtained by adding the uniquely fixed Goldstone-matching contribution implied by the broken-theory Ward identity (cf. Sec. 3). Concretely, the GI completion modifies only the longitudinal entries ($\lambda_{2}=0_{*}$), while leaving the transverse helicity amplitudes unchanged; the additional terms are proportional to the Goldstone-fermion couplings
@@
and a channel-dependent coefficient $c_{G}$ (encoding the normalisation/convention for the Goldstone insertion in the matching). At the amplitude level, the GI completion can be formulated as:
$\widehat{H}^{(\mathrm{GI})}_{\lambda_{0}\lambda_{1}0_{*}}=\widehat{H}^{(\mathrm{SL})}_{\lambda_{0}\lambda_{1}0_{*}}+c_{G}\widehat{H}^{(\mathrm{G_{V}})}_{\lambda_{0}\lambda_{1}0_{*}}$, (30)
with $\widehat{H}^{(\mathrm{G_{V}})}_{\lambda_{0}\lambda_{1}}$ built from the scalar (Goldstone) Yukawa current, i.e. an operator of the type shown in Eq. (29).
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{SL}}_{\lambda_{0}\lambda_{1}\lambda_{2}}=\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{SL}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$ $\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
Table 1: Non-vanishing reduced helicity amplitudes for $q\to q^{\prime}V$ in the SL and GI prescriptions. Entries common to both prescriptions are grouped in the upper block, while the lower block collects the longitudinal ($0_{*}$) components, for which the GI completion adds Goldstone-matching terms proportional to $c_{G}$ and $(y_{L},y_{R})$.
Using the decomposition $F^{(\mathrm{X})}=F_{T}+F_{L}^{(\mathrm{X})}$ and performing the helicity sums implied by the shower construction, weighted by the diagonal entries of the spin-density matrix $\rho$, one obtains the following closed forms for the transverse contribution and for the longitudinal contribution in the SL and GI prescriptions. Writing the kernels as functions of $(z,\tilde{q}^{2})$ (with $\tilde{q}$ the evolution variable of Sec. 2 and with $p_{\perp}^{2}$ eliminated using Eq. (7)), the transverse part is common to both schemes,
@@
whereas the SL contribution reads
$F_{L;\rm{SL}}^{q\to q^{\prime}V}={m_{2}^{2}\over\tilde{q}^{2}(1-z)^{3}}\,\big(g_{R}^{2}\;\rho_{++}+g_{L}^{2}\;\rho_{--}\big),$ (32)
@@
Equations (31)-(33) make explicit the structural point emphasised in Sec. 3: the SL and GI prescriptions share an identical transverse sector, while their longitudinal sectors differ by terms controlled by the Goldstone-matching, including both interference ($\propto c_{G}$) and pure Goldstone ($\propto c_{G}^{2}$) contributions, which are parametrically suppressed for light fermions in Standard Model applications but become essential whenever symmetry-breaking effects and Yukawa insertions compete with the naive $p^{\mu}/M_{V}$ subtraction.